OpenStudy (anonymous):

John inflates 2 balloons in 5 minutes. Alicia ties 5 balloon animals in 12 minutes. They both work at a steady rate. If John maintained this constant rate, how many balloons could he inflate in one hour? If Alicia maintained this constant rate, how many balloon animals could she tie in one hour?

OpenStudy (swissgirl):

Blow 2 ballons in 5 minutes = \( \large\frac{2 Balloons}{5 Minutes}=\frac{2}{5}\) Tie 5 balloon animals in 12 minutes = \( \large \frac{2 Balloons}{12 minutes}=\frac{2}{12}\) So basically We havta blow the balloons and Tie them one hour =60 minutes So we convert the denominator to 60 \( \large\frac{2 Balloons}{5 Minutes}=\frac{2}{5}*\frac{12}{12}=\frac{24}{60}\) So in 1 hour you can blow up 24 Balloons Next we gotta tie the balloons \( \large \frac{2 Balloons}{12 minutes}=\frac{2}{12}*\frac{5}{5}=\frac{10}{60}\) So in one hour you can tie 10 balloons